# A modified Hegselmann–Krause model for interacting voters and political parties

### v2

@@ 2. A modified Hegselmann–Krause model incorporating party dynamics
The noisy Hegselmann–Krause model (1) only models the dynamics of individual voters and is not designed to model the dynamics of a political landscape involving political parties. Therefore, it cannot be used to serve as a model for elections. We will extend the classical Hegselmann–Krause model to couple the preferences and dynamics in the $d$-dimensional opinion space of $N_{v}$ individual voters $v_{i}\in\mathbb{R}^{d}$ and a finite number $N_{p}$ of parties $p_{\alpha}\in\mathbb{R}^{d}$. We make the reasonable assumption that voters will vote for political parties that share similar opinions to them and that both voters and parties, share the same $d$-dimensional opinion space [51]. We make the following assumptions about the interactions between voters and parties: voters are affected by their interactions with other voters and also with parties. In particular, voters attract one another according to the right-hand side drift term of the Hegselmann–Krause model (1) which we denote as $F_{vv}$. The effect is that voters move towards the mean of the voter opinion within a $d$-dimensional sphere with radius $R_{vv}$ centred on the voter. Voters are further attracted by parties which are sufficiently close to them in opinion space [30, 32]. The associated force can be thought of as a leadership effect and will be denoted as $F_{vp}$ [33, 28]. Voters who have a similar opinion to a political party will shape their beliefs to conform to the leaders with whom they are similar. Conversely, parties are affected by the voters and by other parties [31, 38]. It is reasonable to assume that political parties aim to maximise the number of votes they receive. A natural way of achieving this is to minimise the distance between them and their potential voters. However, parties also have “identities” - for example, some are considered left-wing and others centrist - which means that the voters they seek to be close to (or best represent) are within a certain distance $R_{vp}$ in opinion space. We denote the associated attracting force by $F_{pv}$. Finally, it is reasonable to assume that political parties want to differentiate themselves from each other [52]. Why vote for one party if there is another party that is exactly the same? Hence, we assume a repulsive interactive force $F_{pp}$ between political parties. Figure 1 is a schematic of these four interactions.
@@
Here we have chosen force strengths and interaction radii to be equal for all agents. In reality, some parties might be more “charismatic” to certain voters, and similarly, some voters might possess especially strong attractiveness to others (such as between family members). Moreover, some voters might be more open to different opinions and so operate with larger interaction radii than others. To include such agent-specific parameters the parameters in our model (3)–(4) one can make the force strengths and interaction radii voter and party dependent. For example, we could allow for a force strength $\mu^{\alpha}_{{pv}_{i}}$ modelling the leadership effect of party-$\alpha$ on voter-$i$. For simplicity, we do not consider such effects in this work.
@@ 3. Political scenarios in the modified Hegselmann–Krause model
The size of a cluster $\delta_{\rm{cl}}$ is defined as the average diameter of a collection of agents that are not interacting with any other agents, allowing (at least temporarily) for a coherent structure. Heuristically, it is clear that increasing the noise will increase the cluster size, while increasing the force strengths will result in a more tightly packed cluster. We will now find an approximation for the cluster size $\delta_{\rm{cl}}$, in the case where the interaction radii $R_{vv}$ and $R_{pv}$ cover the cluster. Assume for simplicity, that a voter cluster of $N_{v}^{(c)}$ voters is centred around the origin. Further, assume that there are $N_{p}^{(c)}$ parties interacting with the voter cluster and that they are also centred around the origin. Finally, assume that no other agents are interacting with the cluster. With these assumptions the forces acting on the voters become linear with
@@
Yet another important political scenario occurs when parties leave a region in opinion space unoccupied and this region is larger than the interaction radius $R_{pv}$ of any party. In this case voters can occupy this region, attracted by voter-voter interactions only. In this case the voter dynamics degenerates to the original noisy Hegselmann–Krause model (1) and a voter-only cluster forms of size $\delta_{\rm{cl}}\approx 4\sigma_{v}/\sqrt{2\mu_{vv}N_{v}^{(c)}/N_{v}}$ provided that $R_{vv}>\delta_{\rm{cl}}/2$ [48] (cf (14) and (13) with $N_{p}^{(c)}=N_{p}=0$). In this scenario voters may be viewed as disaffected voters who feel unrepresented by political parties, and whose support the parties are not interested in earning. In trying to attract those politically far away voters parties may risk loosing voters that are more aligned with a parties current views. Figure 2(e) shows an example with disaffected voters. Here the same parameters are used as for the party-base clusters in Figure 2(a) but with different initial conditions for the parties allowing for unoccupied political space. The cluster of disaffected voters depicted in Figure 2(e) contains $N_{v}^{{c}}\approx 370$ voters. The estimated cluster size is hence $\delta_{\rm{cl}}\gtrapprox 4\sigma_{v}/\sqrt{2\mu_{vv}0.37}\approx 0.19$ which is consistent with the observed cluster size seen in Figure 2(e) of $0.21$.
Lastly, we report on a scenario in which voters abandon their party-base and aggregate to a single cluster. This occurs when $R_{vv}>R_{pv}$ and $\mu_{vv}\gtrapprox\mu_{pv}$. In this case the attraction between voters is more prominent and the political leadership effect of the parties becomes irrelevant to the behaviour of the voters. Voters collapse to a single state, typically around a single party. This collapsed state we coin in the following voter consensus. Note that this is achieved with all parties having the same force strength $\mu_{pv}$ and without one party having a significantly higher attraction force than the other parties. This may be viewed as voter consensus because the effect occurs when the force between voters becomes more prominent and the political leadership effect of the parties becomes irrelevant to the behaviour of the voters, which collapse to a single state, typically around a single party. An example of such a voter consensus behaviour is seen in Figure 2(f) where a single party-base cluster forms around party $p_{1}$. We remark that if the condition for swing voters (15) is satisfied, the final state will be a swing voter cluster. We have presented here results for a traditional two-party system. The same scenarios also occur in multi-party systems. We present in the Appendix an example for a three-party system.
@@
Figure 2. Prototypical political scenarios for a two-party system. At time $t=0$, $N_{v}=1,000$ voters are distributed uniformly across $[0,1]$ and parties are initially at $p_{1}(0)=0.25$ and $p_{2}(0)=0.75$, except for Figure 2(e). The strengths of the voter forces are $\mu_{vv}=\mu_{pv}=1$. The strengths of the party forces are $\mu_{vp}=0.03$ and $\mu_{pp}=0.01$. The interaction radii are $R_{vv}=0.15$, $R_{pv}=0.1$, $R_{vp}=0.3$ and $R_{vp}=0.2$. The noise strengths are $\sigma_{v}=0.04$ and $\sigma_{p}=0.005$.
@@ 4. Consensus in the modified Hegselmann–Krause model
Figure 4. Evolution of the modified Hegselmann–Krause model (3)–(4) in a $d=1$-dimensional opinion space for $500$ voters (blue) and $4$ parties (red). Initially voters are distributed uniformly across $[0,1]$ and parties are initially at $p_{1}(0)=0.2$, $p_{2}(0)=0.35$, $p_{3}(0)=0.6$ and $p_{4}(0)=0.78$. Parameters are $R_{vv}=0.05$, $R_{pv}=0.1$, $R_{vp}=0.4$, $R_{pp}=0.05$ and $\mu_{vv}=0.5$, $\mu_{pv}=0.8$, $\mu_{vp}=0.02$, $\mu_{pp}=0.02$. The noise strengths are $\sigma_{v}=0.02$ and $\sigma_{p}=0.002$.
@@
(17) $\hat{D}=1-\frac{k_{d}}{3}\left(\frac{1}{N_{v}^{2}}\sum_{v_{i},v_{j}}||v_{i}-v_{j}||+\frac{1}{N_{v}N_{p}}\sum_{v_{i},p_{\alpha}}||v_{i}-p_{\alpha}||+\frac{1}{N_{p}^{2}}\sum_{p_{\alpha},p_{\beta}}||p_{\alpha}-p_{\beta}||\right),$
@@
is the expected distance between two random draws from a uniform distribution on the hypercube $[0,1]^{d}$ with periodic boundary conditions. The parameter $\hat{D}$ is not a strict order parameter in the sense of statistical mechanics but instead a diagnostics measuring in how far the distribution of voters and parties differs from the maximally disordered state of uniformly distributed voters. The proposed consensus diagnostic satisfies $0\leq\hat{D}\leq 1$ with equality given only by uniformity and unanimous consensus, respectively.
@@ 5. Criterion for consensus in the noiseless modified Hegselmann–Krause model
(20) $R>E(0),$
where $R={\rm{min}}(R_{vv},R_{vp},R_{pv},R_{pp})$ is the smallest of all interaction radii and $E(0)$ is the initial value of
@@
which is the convex hull covering the voters and parties with
@@ 6. The mean-field limit and phase transitions
The original noisy Hegselmann–Krause model undergoes a phase transition as the noise strength increases [48] whereby instead of the formation of clusters, voters approach a uniform distribution. In this section, we show that the inclusion of parties does not effect the presence of the phase transition, and we derive an analytical approximation of the critical noise strength, $\sigma_{c}$. This transition from an ordered state of voters to a uniform distribution of voters occurs when $\sigma_{v}>\sigma_{c}$. While large noise is unrealistic – as voters would change their opinions in a random walk – this exercise allows us to study the effect of parties on consensus formation. Perhaps unsurprisingly, we show that parties enhance the formation of voter clusters.
@@ 6.1. One-dimensional case $d=1$
Figure 10 shows a phase diagram obtained from a long simulation of the stationary-party model (32) in $(\sigma_{v},R_{vv})$-space. We consider $2,000$ initially uniformly on $[0,1]$ distributed voters and stationary parties, $p_{1}=0.86,p_{2}=0.53$ and $p_{3}=0.2$. We simulated until time $t=500$ with $\Delta t=0.1$. We tested for statistical equilibrium of the voter dynamics using an Augmented Dickey-Fuller (ADF) test [64], testing for stationarity over the last $20\%$ of the simulation. A phase transition is clearly seen, quantified by the consensus diagnostic $\hat{D}$. To best visualize the phase transition and the departure from uniformity, which for $2,000$ voters uniformly sampled across $[0,1]$ yields $\hat{D}=0.037$, we employ a colour map with the colour transition occurring at $0.039$. The phase transition is well approximated by our approximation (47) for small values of $R_{vv}\lessapprox 0.15$, consistent with the approximation of $R_{vv}\ll 1$ we made to derive (47). The orange region of high level consensus in the phase diagram Figure 10 occurring for $R_{vv}>0.12$ and $\sigma>0.03$ is reached via a transition from a state of three party-base clusters to a single swing voter cluster of smaller size akin to the voter consensus scenario discussed in Section 3 (not shown).
@@ 7. Discussion
The model typically exhibits clusters of voters around individual parties. These clusters form what is known as the party-base of a party. The party-base represents an important feature of politics as it is a core group of voters who consistently support a specific political party [65]. The model further demonstrates the emergence of swing voters as a cluster of voters situated in opinion space between two parties. The parties then compete for the preference of the voters mediated by the repulsive force between them. The distances in opinion space between a voter from the swing cluster and each of the two parties are similar, so small changes in relative party positions can change which party is closest to a particular voter, and hence determines what party they would vote for. The model further supports clusters of disaffected voters that are not aligned with any political party. Such clusters of disaffected voters are generated when political parties evolve into regions in opinion space with large voter mass potentially leaving behind disaffected voters which do not experience any attracting force to the party if their distance is sufficiently large. This latter phenomenon is of increasing importance in modern political science [66]. More extreme political scenarios, such as a sudden collapse of voters to a single party, can be found in the modified Hegselmann–Krause model. We presented results in $1$-dimension and $2$-dimensions that exhibit these behaviours and note that these models tend to approach consensus after a sufficiently long period of time. We remark that in higher dimensions the recurrence of a random walker is not guaranteed [67] and so the merging of all clusters into consensus is also not guaranteed.
@@
While the model assumes that the evolution of voter opinion is due to the relative positions in opinion space, the reality is far more complex. Media consumption and lack of information play an important role in contributing to voters’ decisions [68], which is not covered by the model. Another limitation of the model is the assumption that all parties and all voters have the same interaction radii $R_{pp}$ and $R_{vv}$, respectively, and exert the same force on the other agents. Political parties are clearly not all equal - it is conceivable that some parties exert stronger attractive forces on voters for example, due to political charisma or effective advertising campaigns. Similarly, some voters may be more open to the thoughts and opinions of others, so their interaction radii might be larger than those of other voters. To include such agent-specific force strengths and interaction radii is planned for further research.
@@
Another aspect of political dynamics not examined here is that some political issues may be more controversial than others. While we chose to examine the problem of a cubic domain $[0,1]^{d}$ in this paper, it may be that some issues require more distance to be traversed than others in order for consensus to be achieved. For example, a rectangular domain $[0,L_{1}]\times[0,L_{2}],$ with $L_{2}\gg L_{1}$ would mean that convergence is easier along the smaller dimension because there is less space to traverse. Investigating whether altering the domain significantly alters the model behaviour is left for future work.

### v3

@@ 2. A modified Hegselmann–Krause model incorporating party dynamics
The noisy Hegselmann–Krause model (1) only models the dynamics of individual voters and is not designed to model the dynamics of a political landscape involving political parties. Therefore, it cannot be used to serve as a model for elections. We will extend the classical Hegselmann–Krause model to couple the preferences and dynamics in the $d$-dimensional opinion space of $N_{v}$ individual voters $v_{i}\in\mathbb{R}^{d}$ and a finite number $N_{p}$ of parties $p_{\alpha}\in\mathbb{R}^{d}$. We make the reasonable assumption that voters will vote for political parties that share similar opinions to them and that both voters and parties, share the same $d$-dimensional opinion space [51]. We make the following assumptions about the interactions between voters and parties: voters are affected by their interactions with other voters and also with parties. In particular, voters attract one another according to the right-hand side drift term of the Hegselmann–Krause model (1) which we denote as $F_{vv}$. The effect is that voters move towards the mean of the voter opinion within a $d$-dimensional sphere with radius $R_{vv}$ centred on the voter. Voters are further attracted by parties which are sufficiently close to them in opinion space [30, 32]. The associated force can be thought of as a leadership effect and will be denoted as $F_{pv}$ [33, 28]. Voters who have a similar opinion to a political party will shape their beliefs to conform to the leaders with whom they are similar. Conversely, parties are affected by the voters and by other parties [31, 38]. It is reasonable to assume that political parties aim to maximise the number of votes they receive. A natural way of achieving this is to minimise the distance between them and their potential voters. However, parties also have “identities” - for example, some are considered left-wing and others centrist - which means that the voters they seek to be close to (or best represent) are within a certain distance $R_{vp}$ in opinion space. The smaller the value of $R_{vp}$ the less a party’s view is influenced by far-away voters. We denote the associated attracting force between parties and voters by $F_{vp}$. Finally, it is reasonable to assume that political parties want to differentiate themselves from each other [52]. Why vote for one party if there is another party that is exactly the same? Hence, we assume a repulsive interactive force $F_{pp}$ between political parties. Figure 1 is a schematic of these four interactions.
@@
Here we have chosen force strengths and interaction radii to be equal for all agents. In reality, some parties might be more “charismatic” to certain voters, and similarly, some voters might possess especially strong attractiveness to others (such as between family members). Moreover, some voters might be more open to different opinions and so operate with larger interaction radii than others. To include such agent-specific parameters in our model (3)–(4) one can make the force strengths and interaction radii voter and party dependent. For example, we could allow for a force strength $\mu^{\alpha}_{{pv}_{i}}$ modelling the leadership effect of party-$\alpha$ on voter-$i$. For simplicity, we do not consider such effects in this work.
@@ 3. Political scenarios in the modified Hegselmann–Krause model
The size of a cluster $\delta_{\rm{cl}}$ is defined as the average diameter of a collection of agents that are not interacting with any other agents, allowing (at least temporarily) for a coherent structure. Heuristically, it is clear that increasing the noise will increase the cluster size, while increasing the force strengths will result in a more tightly packed cluster. We will now find an approximation for the cluster size $\delta_{\rm{cl}}$, in the case where the interaction radii $R_{vv}$ and $R_{pv}$ cover the cluster. Suppose that a voter cluster of $N_{v}^{(c)}$ voters is centred around $w\in[0,1]$. Further, assume that there are $N_{p}^{(c)}$ parties interacting with the voter cluster and that they are also centred around the same $w$. As the dynamics depends only on the differences between the positions of the relative agents, we shift, for simplicity, the domain such that the voter cluster is centred at $w=0$. Finally, assume that no other agents are interacting with the cluster. With these assumptions the forces acting on the voters become linear with
@@
Yet another important political scenario occurs when parties leave a region in opinion space unoccupied and this region is larger than the interaction radius $R_{pv}$ of any party. In this case voters can occupy this region, attracted by voter-voter interactions only. The voter dynamics degenerates to the original noisy Hegselmann–Krause model (1) and a voter-only cluster forms of size $\delta_{\rm{cl}}\approx 4\sigma_{v}/\sqrt{2\mu_{vv}N_{v}^{(c)}/N_{v}}$ provided that $R_{vv}>\delta_{\rm{cl}}/2$ [48] (cf (14) and (13) with $N_{p}^{(c)}=N_{p}=0$). In this scenario voters may be viewed as disaffected voters who feel unrepresented by political parties, and whose support the parties are not interested in earning. In trying to attract those politically far away voters parties may risk losing voters that are more aligned with a party’s current views. Figure 2(e) shows an example with disaffected voters. Here the same parameters are used as for the party-base clusters in Figure 2(a) but with different initial conditions for the parties allowing for unoccupied political space. The cluster of disaffected voters depicted in Figure 2(e) contains $N_{v}^{{c}}\approx 370$ voters. The estimated cluster size is hence $\delta_{\rm{cl}}\gtrapprox 4\sigma_{v}/\sqrt{2\mu_{vv}0.37}\approx 0.19$ which is consistent with the observed cluster size seen in Figure 2(e) of $0.21$.
Lastly, we report on a scenario in which voters abandon their party-base and aggregate to a single cluster. This occurs when $R_{vv}>R_{pv}$ and $\mu_{vv}\gtrapprox\mu_{pv}$. In this case the attraction between voters is more prominent and the political leadership effect of the parties becomes irrelevant to the behaviour of the voters. Voters collapse to a single state, typically around a single party. We coin this collapsed state voter consensus. Note that this is achieved with all parties having the same force strength $\mu_{pv}$ and without one party having a significantly higher attraction force than the other parties. This may be viewed as voter consensus because the effect occurs when the force between voters becomes more prominent and the political leadership effect of the parties becomes irrelevant to the behaviour of the voters, which collapse to a single state, typically around a single party. An example of such a voter consensus behaviour is seen in Figure 2(f) where a single party-base cluster forms around party $p_{1}$. We remark that if the condition for swing voters (15) is satisfied, the final state will be a swing voter cluster. We have presented here results for a traditional two-party system. The same scenarios also occur in multi-party systems. We present in the Appendix an example for a three-party system.
@@
Figure 2. Prototypical political scenarios for a two-party system. At time $t=0$, $N_{v}=1,000$ voters are distributed uniformly across $[0,1]$ and parties are initially at $p_{1}(0)=0.25$ and $p_{2}(0)=0.75$, except for Figure 2(e). The strengths of the voter forces are $\mu_{vv}=\mu_{pv}=1$. The strengths of the party forces are $\mu_{vp}=0.03$ and $\mu_{pp}=0.01$. The interaction radii are $R_{vv}=0.15$, $R_{pv}=0.1$, $R_{vp}=0.3$ and $R_{pp}=0.1$. The noise strengths are $\sigma_{v}=0.04$ and $\sigma_{p}=0.005$.
@@ 4. Consensus in the modified Hegselmann–Krause model
Figure 4. Evolution of the modified Hegselmann–Krause model (3)–(4) in a $d=1$-dimensional opinion space for $500$ voters (blue) and $4$ parties (red). Initially voters are distributed uniformly across $[0,1]$ and parties are initially at $p_{1}(0)=0.2$, $p_{2}(0)=0.35$, $p_{3}(0)=0.6$ and $p_{4}(0)=0.78$.
Parameters are $R_{vv}=0.05$, $R_{pv}=0.1$, $R_{vp}=0.4$, $R_{pp}=0.05$ and $\mu_{vv}=0.5$, $\mu_{pv}=0.8$, $\mu_{vp}=0.02$, $\mu_{pp}=0.02$. The noise strengths are $\sigma_{v}=0.02$ and $\sigma_{p}=0.002$.
@@
(17) $\hat{D}=1-\frac{1}{3k_{d}}\left(\frac{1}{N_{v}^{2}}\sum_{v_{i},v_{j}}||v_{i}-v_{j}||+\frac{1}{N_{v}N_{p}}\sum_{v_{i},p_{\alpha}}||v_{i}-p_{\alpha}||+\frac{1}{N_{p}^{2}}\sum_{p_{\alpha},p_{\beta}}||p_{\alpha}-p_{\beta}||\right),$
@@
is the expected distance between two random draws from a uniform distribution on the hypercube $[0,1]^{d}$ with periodic boundary conditions. The parameter $\hat{D}$ is not a strict order parameter in the sense of statistical mechanics but instead a diagnostics measuring how far the distribution of voters and parties differs from the maximally disordered state of uniformly distributed voters. The proposed consensus diagnostic satisfies $0\leq\hat{D}\leq 1$ with equality given only by uniformity and unanimous consensus, respectively.
@@ 5. Criterion for consensus in the noiseless modified Hegselmann–Krause model
(20) $R>\frac{1}{2}\rm{diam}\left(E(0)\right),$
where $R={\rm{min}}(R_{vv},R_{vp},R_{pv},R_{pp})$ is the smallest of all interaction radii and $\rm{diam}(E(0))$ is the diameter of the set
@@
at initial time $t=0$. The set $E(t)$ is the convex hull covering the voters and parties with
@@ 6. The mean-field limit and phase transitions
The original noisy Hegselmann–Krause model undergoes a phase transition as the noise strength increases [48] whereby instead of the formation of clusters, voters approach a uniform distribution. In this section, we show that the inclusion of parties does not affect the presence of the phase transition, and we derive an analytical approximation of the critical noise strength, $\sigma_{c}$. This transition from an ordered state of voters to a uniform distribution of voters occurs when $\sigma_{v}>\sigma_{c}$. While large noise is unrealistic – as voters would change their opinions in a random walk – this exercise allows us to study the effect of parties on consensus formation. Perhaps unsurprisingly, we show that parties enhance the formation of voter clusters.
@@ 6.1. One-dimensional case $d=1$
Figure 10 shows a phase diagram obtained from a long simulation of the stationary-party model (32) in $(\sigma_{v},R_{vv})$-space. We consider $2,000$ voters that are initially distributed uniformly on $[0,1]$ and stationary parties, $p_{1}=0.86,p_{2}=0.53$ and $p_{3}=0.2$. We simulated until time $t=500$ with $\Delta t=0.1$. We tested for statistical equilibrium of the voter dynamics using an Augmented Dickey-Fuller (ADF) test [64], testing for stationarity over the last $20\%$ of the simulation. A phase transition is clearly seen, quantified by the consensus diagnostic $\hat{D}$. To best visualize the phase transition and the departure from uniformity, which for $2,000$ voters uniformly sampled across $[0,1]$ yields $\hat{D}=0.037$, we employ a colour map with the colour transition occurring at $0.039$. The slightly larger value of $\hat{D}=0.039$ was chosen to allow for the detection of sufficiently large deviations from uniformity. The phase transition is well approximated by our approximation (47) for small values of $R_{vv}\lessapprox 0.15$, consistent with the approximation of $R_{vv}\ll 1$ we made to derive (47). Note that the transition in our finite-size system with $2,000$ voters is gradual rather than abrupt as suggested by the mean-field theory. The orange region of high level consensus in the phase diagram in Figure 10 occurring for $R_{vv}>0.12$ and $\sigma>0.03$ is reached via a transition from a state of three party-base clusters to a single swing voter cluster of smaller size akin to the voter consensus scenario discussed in Section 3 (not shown).
@@ 7. Discussion
The model typically exhibits clusters of voters around individual parties. These clusters form what is known as the party-base of a party. The party-base represents an important feature of politics as it is a core group of voters who consistently support a specific political party [65]. The model further demonstrates the emergence of swing voters as a cluster of voters situated in opinion space between two parties. The parties then compete for the preference of the voters mediated by the repulsive force between them. The distances in opinion space between a voter from the swing cluster and each of the two parties are similar, so small changes in relative party positions can change which party is closest to a particular voter, and hence determines what party they would vote for. The model further supports clusters of disaffected voters that are not aligned with any political party. Such clusters of disaffected voters are generated when political parties evolve into regions in opinion space with large voter mass potentially leaving behind disaffected voters which do not experience any attracting force to the party if their distance is sufficiently large. This latter phenomenon is of increasing importance in modern political science [66]. More extreme political scenarios, such as a sudden collapse of voters to a single party, can be found in the modified Hegselmann–Krause model. These scenarios occurred in our model in bounded domains of dimensions $d=1$ and $d=2$. In our simulations, the model approached consensus after a sufficiently long period of time. We remark that in unbounded domains and opinion spaces of dimension $d\geq 3$ recurrence of a random walker is not guaranteed [67] and merging of clusters into unanimous consensus is unlikely.
@@
While the model assumes that the evolution of voter opinion is due to the relative positions in opinion space, the reality is far more complex. Media consumption and lack of information play an important role in contributing to voters’ decisions [68], which is not covered by the model. Another limitation of the model is the assumption that all parties and all voters have the same interaction radii $R_{pp}$ and $R_{vv}$, respectively, and exert the same force on the other agents. Political parties are clearly not all equal - it is conceivable that some parties exert stronger attractive forces on voters for example, due to political charisma or effective advertising campaigns. Similarly, some voters may be more open to the thoughts and opinions of others, so their interaction radii might be larger than those of other voters. In our model parties can freely meander through opinion space, either due to their own stochastic driving force or due to attraction to far away voter clusters (for a sufficiently large interaction radius $R_{vp}$). To ensure that parties remain in some bounded region in opinion space one may include memory in the model, or apply party specific boundary conditions. Inclusion of such agent-specific force strengths and interaction radii is planned for further research.
@@
Another aspect of political dynamics not examined here is that some political issues may be more controversial than others. While we chose to examine the problem of a domain $[0,1]^{d}$ in this paper, it may be that some issues require more distance to be traversed than others in order for consensus to be achieved. For example, a rectangular domain $[0,L_{1}]\times[0,L_{2}],$ with $L_{2}\gg L_{1}$ would mean that convergence is easier along the smaller dimension because there is less space to traverse. Investigating whether altering the domain significantly alters the model behaviour is left for future work.